OpenStudy (mathmath333):

Let p be a prime number such that 17p + 1 is a perfect square. How many prime numbers satisfy this condition?

OpenStudy (mathmath333):

i hope there is only one that is \(\large 19\)

ganeshie8 (ganeshie8):

start with \[17p+1 = n^2\]

ganeshie8 (ganeshie8):

send that \(1\) to other side and factor..

OpenStudy (mathmath333):

\(17p=n^2-1\\ 17p=(n-1)(n+1)\\ \)

ganeshie8 (ganeshie8):

Yes since \(17\) and \(p\) are primes, we they must equal the factors on right hand side in some order

ganeshie8 (ganeshie8):

\[17=n-1, ~~p=n+1\] or \[17=n+1, ~~p=n-1\]

OpenStudy (mathmath333):

\(17=n-1,p=n+1\) \(n=18,p=19\) \(17=n+1,p=n-1\) \(n=16,p=15\)

ganeshie8 (ganeshie8):

yes \(15\) is not a prime so the only solution is \(p=19\)

OpenStudy (mathmath333):

if i use the fact the every prime greater than 3 can be of form \(6n\pm1\) and if i substitute it in the equation and then apply \((mod~~ 6)\) can i prove that

ganeshie8 (ganeshie8):

i feel that gives a more complicated expression..

OpenStudy (mathmath333):

ok i think that will not work because the the converse of \(6n\pm1\) is not true.

ganeshie8 (ganeshie8):

good point